A florist has two arrangements offered at special prices. The first Arrangement consists of 5 lilies and 1 rose and costs $19.75. The second arrangement consists of 6 lilies and 3 roses and costs $27.75. What are the costs of each lily and each rose?

Answer by stanbon(75887)

A florist has two arrangements offered at special prices. The first Arrangement consists of 5 lilies and 1 rose and costs $19.75. The second arrangement consists of 6 lilies and 3 roses and costs $27.75. What are the costs of each lily and each rose?
------------------
Equations:
5L + R = 19.75
6L + 3R = 27.75
--------
Modify for elimination::
15L + 3R = 59.25
6L + 3R = 27.75
-----
Subtract and solve for "L"::
9L = 31.50
L = $3.50 (cost of one Lily)
----
Solve for "R":
5L + R = 19.75
17.50 + R = 19.75
R = $2.50 (cost of one rose)
---------------
Cheers,
Stan H.
--------------